OpenStudy (anonymous):

it looks long but im just explaining what i did and can someone please tell me where i went wrong??? PLEASE? ok, a sector of a circle of radius 6 cm has an angle of 1.6 radians find the area for the sector hence find the area of the segment The problem is that i worked out the area which is 28.8 of the sector i worked out the triangle and took it away from the area of the sector and ended up with 10.8 which is right but then here's the problem how do you use the formula of the area of the sector? it 1/2x6^2xsin(1.6)=0.50258... but that not what i got above i got 10.8 above :S

OpenStudy (amistre64):

the area of a sector is the same as that of a circle; you just replace the angle measurement... 2pi r^2 ------- = area circle 2 angle r^2 -------- = area any part of a circle 2

OpenStudy (amistre64):

1.8pi 6^2 ------- = .9pi(36) = 32.4 pi 2

OpenStudy (amistre64):

ack!!...forget the pi part lol

OpenStudy (amistre64):

32.4

OpenStudy (anonymous):

so is the formula wrong then?

OpenStudy (amistre64):

dunno; but my inputs are dyslexic... 1.6 (36) ------- = area of sector 2

OpenStudy (amistre64):

.8(36) = 28.8

OpenStudy (anonymous):

but thats a different answer :( so is 10.8 wrong then??

OpenStudy (amistre64):

why are you removing the area of the triangle?

OpenStudy (anonymous):

to get the segment area and i looked at the mark scheme and 10.8 is right but theres a segemnt area formula and that doesnt work :(

OpenStudy (amistre64):

so teh segment will be the arch looking section of the piece of pie right?

OpenStudy (amistre64):

the area of a trianlge is: r^2 sin(angle) ------------ 2

OpenStudy (amistre64):

make sure you have your calculator set to radians tho

OpenStudy (anonymous):

sin?? isnt area of a traingle BxH/2???

OpenStudy (amistre64):

it is; but this isnt the area of a right triangle

OpenStudy (anonymous):

oh :o

OpenStudy (amistre64):

the area of any triangle is: base = radial arm = 6 height = radial arm *sin(a) = 6 sin(1.6) and /2 sooo... 36 sin(1.6) --------- = 18 sin(1.6) 2

OpenStudy (amistre64):

the area of the tri is abt: 17.99

OpenStudy (anonymous):

here is the diagram, if u need it

OpenStudy (amistre64):

28.8 -17.99 ------- 10.81

OpenStudy (amistre64):

or maybe not..... is 1.2 larger than 90?

OpenStudy (anonymous):

1.2 lager than 90??? is this a trick qs?

OpenStudy (amistre64):

1.2 * 180 -------- = 68.75 degrees right? pi

OpenStudy (anonymous):

yes, but now im lost where did you get these number fomr?

OpenStudy (amistre64):

trying to convert radians to degrees; radians normally have a pi in them so you multiply by 180/pi to get the angle.. pi 180 -- * ----- = 90 degrees 2 pi

OpenStudy (amistre64):

12 180 --- * ---- = 216/pi = 68.75 10 pi

OpenStudy (anonymous):

oh :o

OpenStudy (amistre64):

since the angle is acute; we can assume that our method was correct;

OpenStudy (amistre64):

so I get 10.81 whats the answer you get?

OpenStudy (anonymous):

10.8 :)

OpenStudy (amistre64):

then the area of the segment should be 10.8 :)

OpenStudy (anonymous):

thank you soo much for explaining this :) if i could i would have wanted to give you some more medals but i can only give you one :S

OpenStudy (amistre64):

'sok; you enjoy yourself now :) i got class starting soon

OpenStudy (anonymous):

bye bye :)